OpenStudy (anonymous):

4x^3=108 : Find the real number solutions. Please help!!!

OpenStudy (jdoe0001):

\(\bf 4x^3=108\implies x^3=\cfrac{108}{4}\implies x=\sqrt[3]{\cfrac{108}{4}}\)

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

http://mathb.in/12750

OpenStudy (destinymasha):

do you have options? It will make work much faster

OpenStudy (destinymasha):

if not, estimate , and substitute for x

OpenStudy (anonymous):

How?

OpenStudy (destinymasha):

you could divide each side by 4

OpenStudy (destinymasha):

4 divided by 4 = 1 , and 108 divided by 4 = 27. so the equation becomes x^3 = 27

OpenStudy (destinymasha):

then substitute for x

OpenStudy (destinymasha):

then x would equal 3 because 3^3 = 27.

OpenStudy (anonymous):

So the answer's 3?

OpenStudy (destinymasha):

yup

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You sure?

OpenStudy (destinymasha):

http://www.tiger-algebra.com/drill/4x~3=108/